OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 3x. What is the value of (g o f)(3)? 6 x+8 -10 23

OpenStudy (anonymous):

what are you trying to say? @SwoleNation_2016

jimthompson5910 (jim_thompson5910):

dav21 you have the order in reverse and SwoleNation_2016 that's irrelevant and inappropriate, megen505 looks just fine, so stop trying to troll or you may be banned

OpenStudy (anonymous):

You know, she is on here to get help with a math problem. She isn't ugly, whatsoever. Cyberbulling is terrible.

jimthompson5910 (jim_thompson5910):

kbbbb agreed, just too much negativity out there

OpenStudy (anonymous):

woooowww your nice

jimthompson5910 (jim_thompson5910):

SwoleNation_2016 you've been reported knock that off right now

OpenStudy (anonymous):

i dont even know you?

OpenStudy (anonymous):

@megen505 some people just live to be rude and hurt people.

jimthompson5910 (jim_thompson5910):

anyways, back to math (g o f)(x) = g(f(x)) so we need to find g(f(x)) first ------------------------------------------------------- g(x) = 3x g(f(x)) = 3( f(x) ) ... replace each x with f(x) g(f(x)) = 3( -2x + 8 ) ... replace each f(x) on the right side with -2x + 8 g(f(x)) = -6x + 24

jimthompson5910 (jim_thompson5910):

so this means that (g o f)(x) = -6x + 24 now plug in x = 3 (g o f)(x) = -6x + 24 (g o f)(3) = -6*3 + 24 (g o f)(3) = ???

OpenStudy (anonymous):

the answer is 6

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (ranga):

f(x) = -2x + 8 and g(x) = 3x. What is g(f(x)) when x = 3? First solve for f(x) when x = 3. f(3) = (-2)(3) + 8 = 2 g(f(3)) = g(2) = (3)(2) = 6.

OpenStudy (anonymous):

thank you so much @ranga

OpenStudy (ranga):

You are welcome. Don't think too much about troll remarks. Enjoy the rest of your day.